Daily high temperatures in St. Louis for the last week were as follows: 33, 33, 38,...

  1. Daily high temperatures in St. Louis for the last week were as follows: 33, 33, 38, 36, 43, 23, 28 (yesterday).
    1. Forecast the high temperature today, using a 3-day moving average.
    2. Forecast the high temperature today, using a 2-day moving average.

For questions c-e use all the days where a 3-day forecast value is possible.

  1. Calculate the mean absolute deviation based on a 3-day moving average.
  2. Compute the mean squared error for the 3-day moving average.
  3. Calculate the mean absolute percent error for the 3-day moving averag
Answer #1

Forecast = Sigma(Demand in previous N periods) / N

1. 2PMA

FORECAST 8 = (23 + 28) / 2 = 25.5

2. 3PMA

FORECAST 8 = (43 + 23 + 28) / 3 = 31.33


MAD, MSE, AND MAPE

MAD = SIGMA(SUM OF ABSOLUTE DEVIATIONS) / N

MSE = SIGMA(SUM OF DEVIATIONS) ^ 2 / N

MAPE = SIGMA(SUM OF ABS DEV / DEMAND * 100) / N

PERIOD

ACTUAL DEMAND

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

ABS DEV / DEMAND

DEVIATIONS^2

1

33

2

33

3

38

4

36

34.67

36 - 34.67 = 1.33

1.33

0.036944

1.7689

5

43

35.67

43 - 35.67 = 7.33

7.33

0.170465

53.7289

6

23

39

23 - 39 = -16

16

0.695652

256

7

28

34

28 - 34 = -6

6

0.214286

36

SIGMA

-13.34

30.66

1.117347

347.4978

N = 4

MAD = 30.66 / 4 = 7.67

MSE = 347.4978 / 4 = 86.87455 OR 86.88

MAPE = (1.117347 * 100) / 4 = 27.93(rounded)

**PLEASE RATE WITH A THUMBS UP :)